# The Loss of Control Playbook: Degrees, Dynamics, and Preparedness

### v1

@@ 2.B. Focus on Deployment Environment, Affordances, and Permissions
For decision- and policymakers who wish to implement mechanisms and interventions to safeguard against LoC today, uncertainties around capabilities and propensities and their causal impact on LoC, pose a significant hurdle. We therefore propose a complementary approach that can offer actionable levers today, while largely sidestepping uncertainties surrounding capabilities and propensities. Specifically, as opposed to focusing on AI systems’ intrinsic factors (i.e., capabilities and propensities), we propose to focus on extrinsic factors that can raise the overall risk of LoC occurring. In doing so, we propose a framework inspired by the research conducted in Chapter 1, focusing on deployment context, affordances, and permissions. We refer to this framework as the DAP framework.
@@ 2.B.1 Deployment Context
In order to assist decision- and policymakers, we put forward some proposals for what they might wish to consider as high-stakes deployment contexts. We believe that an initial list of high-stakes deployment contexts can be inferred from the literature review in our Chapter 1, and the relevant graph (see Figure 2). [^25] More details can be found in Chapter 1 and in Figure 2. Specifically, in our review of concrete scenarios, we found that there were certain recurring deployment environments that appeared especially high-stakes. For example:
  • Critical national infrastructure. We found two concrete LoC scenarios evolving within critical infrastructure, ranging from large-scale electricity outages (‘Scenario 1’) to catastrophic failures across critical sectors (‘Scenario 4’) [Kalra_GeopoliticsAGI]. This finding aligns with other categorizations of critical environments, as, for example, seen in the USA Patriot Act of 2001 as “systems and assets … so vital to the United States that the incapacity or destruction … would have a debilitating impact on security, national economic security, national public health or safety, or any combination of those matters” [USC42_5195c, ppd21_2013]. [^26] The sectors included in CNI may vary between jurisdictions or legislative purposes. For example, the EU’s cybersecurity NIS-2 directive includes energy, transport, banking, financial market infrastructure, health, drinking water, waste water, digital infrastructure, information and communication technology service management, public administration and space as sectors with “high-criticality” [DirectiveEU2022_2555_NIS2]. The U.S. Cybersecurity and Infrastructure Agency does not include some of these, but includes sectors such as dams, defense industrial base and nuclear reactors, materials, and waste [CISA_Critical_Infrastructure_Sectors].
  • Military. We found two concrete LoC scenarios relating to AI deployment in a military context, involving AI systems being tasked with military coordination in ‘Scenario 3’ in [Kalra_GeopoliticsAGI] or autonomous weapons [ai2027:site]. These findings reflect existing concerns around AI systems becoming instrumental for military decision-making under strategic pressure, which could lead to conflict escalation [GlobalCommissionREAIM2025, Rivera_2024, HoffmanKim2023ReducingRisksAI], as well as around autonomous weapons systems [USDeptState2023PoliticalDeclaration, UNGA_ARES79239_2024, UNGA_ARES78241_2023], which could pose “serious concerns from humanitarian, legal and ethical perspectives” [ICRC2021AutonomousWeaponsPosition]. Moreover, this finding aligns with the fact that military operations are considered to be inherently risky [CJCSI3213_01D_2012]. [^27] An additional example pertains to army operations directly which are considered “inherently dangerous” according to the Risk Management Pamphlet of the Department of the Army [DA_Pam_385_30_2014].
@@ 2.B.3 Permissions
Third, we turn our focus toward permissions. Permissions given to an AI system determine what actions it is authorized to take [stix2025aicloseddoorsprimer, nist_csrc_permission_glossary]. In other words, the term ’permission’ refers to whether an AI system is enabled by its developers to utilize its capabilities through the available affordances. Permissions are closely intertwined with affordances and ought to be considered in tandem to provide either an enabling or a restrictive function. For example, an AI system may have technical access to a social media site through its affordances, for example, to read posts, but may lack the necessary permission to publish posts. Conversely, it may have permission to publish posts, but lack the affordance to access the corresponding website in the first place; in this case, it cannot enact the permission until it also has the affordance.
@@ 2.C Reflections
In this Chapter, we aim to model what happens should AI capabilities, including those relevant to LoC, continue to increase, and should highly advanced future AI systems be deployed in more complex and high-stakes deployment contexts, with broad affordances and permissions. We claim that, under these conditions, society would eventually find itself living with a ‘state of vulnerability’ (see Section 3.A.). We propose the term state of vulnerability to denote a state in which a sufficiently capable future AI system has acquired (through humans or independently) or could independently acquire sufficient access to resources, affordances, and permissions (or means to acquire further access) and sufficient capabilities to cause LoC when a catalyst materializes.
We propose that misalignment and pure malfunction (see Section 3.B) are the catalyst for LoC. In other words, they are the spark igniting an otherwise inactive bomb. Once society is living with a state of vulnerability, we have essentially created an environment akin to sitting on a powder keg. Once ignited through a catalyst, i.e., misalignment and pure malfunction, the situation is likely to devolve into LoC (see Section 3.C.).  [^31] It is plausible that there are multiple pathways by which misalignment and pure malfunction could occur, such as by misuse. The specificities of these pathways are outside of the scope of this report.
More specifically, through our theoretical mapping of plausible future pathways, we speculate that once society is living with a state of vulnerability, the majority of pathways would eventually lead to LoC materializing. At the very least, we propose that it would be impossible to meaningfully assure ex ante that LoC will not materialize. In our mapping, only a small minority of pathways would maintain a ‘safe world,’ i.e., a situation where the deployment of a highly advanced future AI system never devolves into LoC for the entirety of its service period. We therefore conclude that it is imperative for society to either: (i) ensure a state of vulnerability is not reached, which is implausible given economic and strategic pressures (see Section 3.A); to (ii) ensure that any given highly advanced future AI system is not deployed if an immediate materialization of LoC seems realistic, which will be hard or impossible to determine ex ante (see Section 3.C.); or to (iii) develop a sufficient defense-in-depth approach including safeguards, oversight and control mechanisms to hold any given highly advanced future AI system in a perennial state of suspension vis-à-vis LoC materializing.
@@ 3.D Reflections
  • Technical interventions including, among others: rigorous pre-deployment testing suites in accordance with threat models for the deployment context; control measures that constrain an AI system’s effect on the world around the AI system ([greenblatt2024aicontrolimprovingsafety, korbak2025sketchaicontrolsafety]; as well as stringent human and AI-enabled monitoring [ncsc_secure_ai_2023]. [^35] We want to be explicit that: (i) the spectrum of these interventions may change and that interventions should adapt to the most robust future mechanisms available; that (ii) these interventions are not comprehensive; and that (ii) these intervention categories should not be misconstrued as containing independently sufficient features but instead should always be considered holistically.
@@ Acknowledgements
For thoughtful feedback and valuable discussions related to this work, we would like to thank: David Africa, Jacob Arbeid, Benjamin Boudreaux, Axel Højmark, Marius Hobbhahn, Alex Meinke, Sören Mindermann, Jérémy Scheurer, Anna Wang, Teun van der Weij, Caleb Withers, and multiple reviewers who wish to stay anonymous.
@@
In this Appendix, we provide additional context on the assumptions and methodology underpinning our taxonomy of LoC described in Chapter 1 and visually represented in Figure 2. Specifically, we provide additional context on:
@@ 2.1 Literature Reviewed
Carlsmith, Joseph. 2024. ”Is Power-Seeking AI an Existential Risk?” https://arxiv.org/abs/2206.13353. Critch, Andrew, and Stuart Russell. 2023. ”TASRA: a Taxonomy and Analysis of Societal-Scale Risks from AI”. https://arxiv.org/abs/2306.06924.
Barnett, Peter, and Jeremy Gillen. 2024. ”Without Fundamental Advances, Misalignment and Catastrophe Are the Default Outcomes of Training Powerful AI”. https://intelligence.org/wp-content/uploads/2024/12/MisalignmentandCatastrophe.pdf.
@@
Bengio, Yoshua, Michael Cohen, Damiano Fornasiere, Joumana Ghosn, Pietro Greiner, et al. 2025. ”Superintelligent Agents Pose Catastrophic Risks: Can Scientist AI Offer a Safer Path?”. https://arxiv.org/abs/2502.15657.
@@
Shah, Rohin, Vikrant Varma, Ramana Kumar, Mary Phuong, Victoria Krakovna, et al. 2022. ”Goal Misgeneralization: Why Correct Specifications Aren’t Enough For Correct Goals”. https://arxiv.org/abs/2210.01790.
@@
Sotala, Kaj. 2018. ”Disjunctive Scenarios of Catastrophic AI Risk”. https://kajsotala.fi/assets/2018/12/Disjunctivescenarios.pdf.
@@
Kulveit, Jan, Raymond Douglas, Nora Ammann, Deger Turan, David Krueger, et al. 2025. ”Gradual Disempowerment: Systemic Existential Risks from Incremental AI Development”. https://arxiv.org/abs/2501.16946.
@@
Clarke, Sam, Jess Whittlestone. 2022. ”A Survey of the Potential Long-term Impacts of AI.” https://dl.acm.org/doi/pdf/10.1145/3514094.3534131.
@@
Hendrycks, Dan, Eric Schmidt, Alexandr Wang. 2025. ”Superintelligence Strategy: Expert Version.” https://arxiv.org/pdf/2503.05628.
@@
Hendrycks, Dan, Mantas Mazeika, and Thomas Woodside. 2023. ”An Overview of Catastrophic AI Risks”. https://arxiv.org/abs/2306.12001.
@@
Ngo, Richard, Lawrence Chan, Sören Mindermann. 2025. ”The Alignment Problem from a Deep Learning Perspective”. https://arxiv.org/pdf/2209.00626.
@@
Pavel, Barry, Ivana Ke, Gregory Smith, Sophia Brown-Heidenreich, Lea Sabbag, et al. 2025. ”How Artificial General Intelligence Could Affect the Rise and Fall of Nations: Visions for Potential AGI Futures.” https://doi.org/10.7249/RRA3034-2.
@@
Özcan, Bengüsu, Daan Juijn, Jakob Graabak, Sam Bogerd. 2025. ”Advanced AI: Possible futures”. https://cfg.eu/advanced-ai-possible-futures/#take-off.
@@
Bengio, Yoshua. 2024. ”Government Interventions to Avert Future Catastrophic AI Risks”. https://hdsr.mitpress.mit.edu/pub/w974bwb0/release/2.
@@
Hendrycks, Dan. 2023. ”Natural Selection Favors AIs over Humans”. https://arxiv.org/abs/2303.16200.
@@
Shah, Rohin, Alex Irpan, Alexander Matt Turner, Anna Wang, Arthur Conmy et al. 2025. ”An Approach to Technical AGI Safety and Security”. https://arxiv.org/pdf/2504.01849.
@@
Raman, Deepika, Nada Madkour, Evan R. Murphy, Krystal Jackson, Jessica Newman. 2025. ”Intolerable Risk Threshold Recommendations for Artificial Intelligence” https://www.aigl.blog/content/files/2025/04/Intolerable-Risk-Threshold-Recommendations-for-Artificial-Intelligence.pdf.
@@
Somani, Elika, Anjay Friedman, Henry Wu, Marianne Lu, Christopher Byrd, et al. 2025. ”Strengthening Emergency Preparedness and Response for AI Loss of Control Incidents.” https://doi.org/10.7249/RRA3847-1.
@@
Wasil, Akash R., Everett Smith, Corin Katzke, Justin Bullock. 2024. ”AI Emergency Preparedness: Examining the federal government’s ability to detect and respond to AI-related national security threats” https://arxiv.org/pdf/2407.17347.
@@
Salmon, Paul M., Tony Carden, Peter A. Hancock. 2020. ”Putting the humanity into inhuman systems: How human factors and ergonomics can be used to manage the risks associated with artificial general intelligence”. https://doi.org/10.1002/hfm.20883.
@@
Department for Science, Innovation & Technology (DSIT). 2025. ”Future risks of frontier AI (Annex A)”. https://www.gov.uk/government/publications/frontier-ai-capabilities-and-risks-discussion-paper/future-risks-of-frontier-ai-annex-a.
@@
Bernardi, Jamie, Gabriel Mukobi, Hilary Greaves, Lennart Heim, Markus Anderljung. 2025. ”Societal Adaptation to Advanced AI”. https://arxiv.org/pdf/2405.10295.
@@
Drexel, Bill, Caleb Withers. 2024. ”Catalyzing Crisis: A Primer on Artificial Intelligence, Catastrophes, and National Security.” https://s3.us-east-1.amazonaws.com/files.cnas.org/documents/Catastrophic-AI˙TECH-2024˙Final.pdf.
@@
Bengio, Yoshua, Sören Mindermann, Daniel Privitera, Tamay Besiroglu, Rishi Bommasani, et al. 2025. ”International AI Safety Report 2025.” https://internationalaisafetyreport.org/publication/international-ai-safety-report-2025.
@@
Cass-Beggs, Duncan, Stephen Clare, Dawn Dimowo, Zaheed Kara. 2024. ”Framework Convention on Global AI Challenges: Accelerating International Cooperation to Ensure Beneficial, Safe and Inclusive AI.” https://www.cigionline.org/static/documents/AI-challenges˙OW6rTMD.pdf.
@@
Bengio, Yoshua, Geoffrey Hinton, Andrew Yao, Dawn Song, Pieter Abbeel, et al. 2024. ”Managing Extreme AI Risks amid Rapid Progress.” https://arxiv.org/abs/2310.17688.
@@
Barnett, Peter, Aaron Scher, David Abecassis. 2025. ”Technical Requirements for Halting Dangerous AI Activities.” https://arxiv.org/abs/2507.09801.
@@
Park, Peter S., Simon Goldstein, Aidan O’Gara, Michael Chen, Dan Hendrycks. 2023. ”AI Deception: A Survey of Examples, Risks, and Potential Solutions.” https://arxiv.org/abs/2308.14752.
@@
Bales, Adam, William D’Alessandro, Cameron Domenico Kirk-Giannini. 2024. ”Artificial Intelligence: Arguments for Catastrophic Risk.” https://arxiv.org/abs/2401.15487.
@@
Motwani, Sumeet Ramesh, Mikhail Baranchuk, Martin Strohmeier, Vijay Bolina, Philip H.S. Torr, et al. 2025. ”Secret Collusion among AI Agents: Multi-Agent Deception via Steganography.” https://arxiv.org/abs/2402.07510.
@@
Barnett, Peter, Aaron Scher. 2025. ”AI Governance to Avoid Extinction: The Strategic Landscape and Actionable Research Questions.” https://arxiv.org/abs/2505.04592.
@@
Jones, Charles I. 2024. ”The AI Dilemma: Growth versus Existential Risk.” https://web.stanford.edu/˜chadj/existentialrisk.pdf.
@@
Swoboda, Torben, Risto Uuk, Lode Lauwaert, Andrew P. Rebera, Ann-Katrien Oimann, et al. 2025. ”Examining Popular Arguments Against AI Existential Risk: A Philosophical Analysis.” https://arxiv.org/pdf/2501.04064.
@@
Miotti, Andrea. 2025. ”Taking control: Policies to address extinction risks from advanced AI.” https://arxiv.org/pdf/2310.20563.
@@
Phuong, Mary, Roland S. Zimmermann, Ziyue Wang, David Lindner, Victoria Krakovna, et al. 2025. ”Evaluating Frontier Models for Stealth and Situational Awareness.” https://arxiv.org/pdf/2505.01420.
@@
Bucknall, Benjamin S., Shiri Dori-Hacohen. 2022. ”Current and Near-Term AI as a Potential Existential Risk Factor.” https://arxiv.org/abs/2209.10604.
@@
Russell, Stuart. 2022. ”Artificial Intelligence and the Problem of Control.” https://link.springer.com/chapter/10.1007/978-3-030-86144-5˙3.
Li, Xiaojian, Haoyuan Shi, Rongwu Xu, Wei Xu. 2025. ”AI Awareness.” https://arxiv.org/pdf/2504.20084.
@@
Dung, Leonard. 2025. ”Evaluating approaches for reducing catastrophic risks from AI.” https://link.springer.com/article/10.1007/s43681-024-00475-w.
@@
Kingsley, Jeremy, Syedah Ailia Haider, Shreyansh Jain, Adam Green, Jan Copeman, et al. 2024. ”AI landscapes: Exploring future scenarios of AI through to 2030.” https://securesustain.org/wp-content/uploads/2024/05/AI-Landscapes-Exploring-Future-Scenarios-of-AI-Through-to-2030.pdf.
@@
Lynch, Aengus, Benjamin Wright, Caleb Larson, Stuart J. Ritchie, Kevin K. Troy, et al. 2025. ”Agentic Misalignment: How LLMs could be insider threats.” https://www.anthropic.com/research/agentic-misalignment.
@@
Benavoli, Alessio, Alessandro Facchini, Marco Zaffalon. 2025. ”The AI off-switch problem as a signalling game: bounded rationality and incomparability.” https://arxiv.org/abs/2502.06403.

### v2

@@ 2.B. Focus on Deployment Environment, Affordances, and Permissions
For decision- and policymakers who wish to implement mechanisms and interventions to safeguard against LoC today, uncertainties around capabilities and propensities, and their causal impact on LoC, pose a significant hurdle. We therefore propose a complementary approach that can offer actionable levers today, while largely sidestepping uncertainties surrounding capabilities and propensities. Specifically, as opposed to focusing on AI systems’ intrinsic factors (i.e., capabilities and propensities), we propose to focus on extrinsic factors that can raise the overall risk of LoC occurring. In doing so, we propose a framework inspired by the research conducted in Chapter 1, focusing on deployment context, affordances, and permissions. We refer to this framework as the DAP framework.
@@ 2.B.1 Deployment Context
In order to assist decision- and policymakers, we put forward some proposals for what they might wish to consider as high-stakes deployment contexts. We believe that an initial list of high-stakes deployment contexts can be inferred from the literature review in our Chapter 1, and the relevant graph (see footnote 15). [^25] More details can be found in Chapter 1 and in footnote 15. Specifically, in our review of concrete scenarios, we found that there were certain recurring deployment environments that appeared especially high-stakes. For example:
  • Critical national infrastructure. We found two concrete LoC scenarios evolving within critical infrastructure, ranging from large-scale electricity outages (‘Scenario 1’) to catastrophic failures across critical sectors (‘Scenario 4’) in [Kalra_GeopoliticsAGI]. This finding aligns with other categorizations of critical environments, as, for example, seen in the USA Patriot Act of 2001 as “systems and assets … so vital to the United States that the incapacity or destruction … would have a debilitating impact on security, national economic security, national public health or safety, or any combination of those matters” [USC42_5195c, ppd21_2013]. [^26] The sectors included in CNI may vary between jurisdictions or legislative purposes. For example, the EU’s cybersecurity NIS-2 directive includes energy, transport, banking, financial market infrastructure, health, drinking water, waste water, digital infrastructure, information and communication technology service management, public administration and space as sectors with “high-criticality” [DirectiveEU2022_2555_NIS2]. The U.S. Cybersecurity and Infrastructure Agency does not include some of these, but includes sectors such as dams, defense industrial base and nuclear reactors, materials, and waste [CISA_Critical_Infrastructure_Sectors].
  • Military. We found two concrete LoC scenarios relating to AI deployment in a military context, involving AI systems being tasked with military coordination in ‘Scenario 3’ in [Kalra_GeopoliticsAGI] and in [ai2027:site]. These findings reflect existing concerns around AI systems becoming instrumental for military decision-making under strategic pressure, which could lead to conflict escalation [GlobalCommissionREAIM2025, Rivera_2024, HoffmanKim2023ReducingRisksAI], as well as around autonomous weapons systems [USDeptState2023PoliticalDeclaration, UNGA_ARES79239_2024, UNGA_ARES78241_2023], which could pose “serious concerns from humanitarian, legal and ethical perspectives” [ICRC2021AutonomousWeaponsPosition]. Moreover, this finding aligns with the fact that military operations are considered to be inherently risky [CJCSI3213_01D_2012]. [^27] An additional example pertains to army operations directly which are considered “inherently dangerous” according to the Risk Management Pamphlet of the Department of the Army [DA_Pam_385_30_2014].
@@ 2.B.3 Permissions
Third, we turn our focus toward permissions. Permissions given to an AI system determine what actions it is authorized to take [stix2025aicloseddoorsprimer, nist_csrc_permission_glossary]. In other words, the term ‘permission’ refers to whether an AI system is enabled by its developers to utilize its capabilities through the available affordances. Permissions are closely intertwined with affordances and ought to be considered in tandem to provide either an enabling or a restrictive function. For example, an AI system may have technical access to a social media site through its affordances, for example, to read posts, but may lack the necessary permission to publish posts. Conversely, it may have permission to publish posts, but lack the affordance to access the corresponding website in the first place; in this case, it cannot enact the permission until it also has the affordance.
@@ 2.C Reflections
In this Chapter, we aim to model what happens should AI capabilities, including those relevant to LoC, continue to increase, and should highly advanced future AI systems be deployed in more complex and high-stakes deployment contexts, with broad affordances and permissions. We claim that, under these conditions, society would eventually find itself living with a ‘state of vulnerability’ (see Section 3.A). We propose the term state of vulnerability to denote a state in which a sufficiently capable future AI system has acquired (through humans or independently) or could independently acquire sufficient access to resources, affordances, and permissions (or means to acquire further access) and sufficient capabilities to cause LoC when a catalyst materializes.
We propose that misalignment and pure malfunction (see Section 3.B) are the catalyst for LoC. In other words, they are the spark igniting an otherwise inactive bomb. Once society is living with a state of vulnerability, we have essentially created an environment akin to sitting on a powder keg. Once ignited through a catalyst, i.e., misalignment and pure malfunction, the situation is likely to devolve into LoC (see Section 3.C).  [^31] It is plausible that there are multiple pathways by which misalignment and pure malfunction could occur, such as by misuse. The specificities of these pathways are outside of the scope of this report.
More specifically, through our theoretical mapping of plausible future pathways, we speculate that once society is living with a state of vulnerability, the majority of pathways would eventually lead to LoC materializing. At the very least, we propose that it would be impossible to meaningfully assure ex ante that LoC will not materialize. In our mapping, only a small minority of pathways would maintain a ‘safe world,’ i.e., a situation where the deployment of a highly advanced future AI system never devolves into LoC for the entirety of its service period. We therefore conclude that it is imperative for society to either: (i) ensure a state of vulnerability is not reached, which is implausible given economic and strategic pressures (see Section 3.A); to (ii) ensure that any given highly advanced future AI system is not deployed if an immediate materialization of LoC seems realistic, which will be hard or impossible to determine ex ante (see Section 3.C); or to (iii) develop a sufficient defense-in-depth approach including safeguards, oversight and control mechanisms to hold any given highly advanced future AI system in a perennial state of suspension vis-à-vis LoC materializing.
@@ 3.D Reflections
  • Technical interventions including, among others: rigorous pre-deployment testing suites in accordance with threat models for the deployment context; control measures that constrain an AI system’s effect on the world around the AI system ([greenblatt2024aicontrolimprovingsafety, korbak2025sketchaicontrolsafety]; as well as stringent human and AI-enabled monitoring [ncsc_secure_ai_2023]. [^35] We want to be explicit that: (i) the spectrum of these interventions may change and that interventions should adapt to the most robust future mechanisms available; that (ii) these interventions are not comprehensive; and that (iii) these intervention categories should not be misconstrued as containing independently sufficient features but instead should always be considered holistically.
@@ Acknowledgements
For thoughtful feedback and valuable discussions related to this work, we would like to thank: David Africa, Jacob Arbeid, Benjamin Boudreaux, Axel Højmark, Marius Hobbhahn, Alex Meinke, Sören Mindermann, Jérémy Scheurer, Anna Wang, Teun van der Weij, Caleb Withers, Mikita Balesni and multiple reviewers who wish to stay anonymous.
@@
In this Appendix, we provide additional context on the assumptions and methodology underpinning our taxonomy of LoC described in Chapter 1 and visually represented in footnote 15. Specifically, we provide additional context on:
@@ 2.1 Literature Reviewed
Carlsmith, Joseph. 2024. “Is Power-Seeking AI an Existential Risk?” https://arxiv.org/abs/2206.13353. Critch, Andrew, and Stuart Russell. 2023. “TASRA: a Taxonomy and Analysis of Societal-Scale Risks from AI”. https://arxiv.org/abs/2306.06924.
Barnett, Peter, and Jeremy Gillen. 2024. “Without Fundamental Advances, Misalignment and Catastrophe Are the Default Outcomes of Training Powerful AI”. https://intelligence.org/wp-content/uploads/2024/12/MisalignmentandCatastrophe.pdf.
@@
Bengio, Yoshua, Michael Cohen, Damiano Fornasiere, Joumana Ghosn, Pietro Greiner, et al. 2025. “Superintelligent Agents Pose Catastrophic Risks: Can Scientist AI Offer a Safer Path?”. https://arxiv.org/abs/2502.15657.
@@
Shah, Rohin, Vikrant Varma, Ramana Kumar, Mary Phuong, Victoria Krakovna, et al. 2022. “Goal Misgeneralization: Why Correct Specifications Aren’t Enough For Correct Goals”. https://arxiv.org/abs/2210.01790.
@@
Sotala, Kaj. 2018. “Disjunctive Scenarios of Catastrophic AI Risk”. https://kajsotala.fi/assets/2018/12/Disjunctivescenarios.pdf.
@@
Kulveit, Jan, Raymond Douglas, Nora Ammann, Deger Turan, David Krueger, et al. 2025. “Gradual Disempowerment: Systemic Existential Risks from Incremental AI Development”. https://arxiv.org/abs/2501.16946.
@@
Clarke, Sam, Jess Whittlestone. 2022. “A Survey of the Potential Long-term Impacts of AI.” https://dl.acm.org/doi/pdf/10.1145/3514094.3534131.
@@
Hendrycks, Dan, Eric Schmidt, Alexandr Wang. 2025. “Superintelligence Strategy: Expert Version.” https://arxiv.org/pdf/2503.05628.
@@
Hendrycks, Dan, Mantas Mazeika, and Thomas Woodside. 2023. “An Overview of Catastrophic AI Risks”. https://arxiv.org/abs/2306.12001.
@@
Ngo, Richard, Lawrence Chan, Sören Mindermann. 2025. “The Alignment Problem from a Deep Learning Perspective”. https://arxiv.org/pdf/2209.00626.
@@
Pavel, Barry, Ivana Ke, Gregory Smith, Sophia Brown-Heidenreich, Lea Sabbag, et al. 2025. “How Artificial General Intelligence Could Affect the Rise and Fall of Nations: Visions for Potential AGI Futures.” https://doi.org/10.7249/RRA3034-2.
@@
Özcan, Bengüsu, Daan Juijn, Jakob Graabak, Sam Bogerd. 2025. “Advanced AI: Possible futures”. https://cfg.eu/advanced-ai-possible-futures/#take-off.
@@
Bengio, Yoshua. 2024. “Government Interventions to Avert Future Catastrophic AI Risks”. https://hdsr.mitpress.mit.edu/pub/w974bwb0/release/2.
@@
Hendrycks, Dan. 2023. “Natural Selection Favors AIs over Humans”. https://arxiv.org/abs/2303.16200.
@@
Shah, Rohin, Alex Irpan, Alexander Matt Turner, Anna Wang, Arthur Conmy et al. 2025. “An Approach to Technical AGI Safety and Security”. https://arxiv.org/pdf/2504.01849.
@@
Raman, Deepika, Nada Madkour, Evan R. Murphy, Krystal Jackson, Jessica Newman. 2025. “Intolerable Risk Threshold Recommendations for Artificial Intelligence” https://www.aigl.blog/content/files/2025/04/Intolerable-Risk-Threshold-Recommendations-for-Artificial-Intelligence.pdf.
@@
Somani, Elika, Anjay Friedman, Henry Wu, Marianne Lu, Christopher Byrd, et al. 2025. “Strengthening Emergency Preparedness and Response for AI Loss of Control Incidents.” https://doi.org/10.7249/RRA3847-1.
@@
Wasil, Akash R., Everett Smith, Corin Katzke, Justin Bullock. 2024. “AI Emergency Preparedness: Examining the federal government’s ability to detect and respond to AI-related national security threats” https://arxiv.org/pdf/2407.17347.
@@
Salmon, Paul M., Tony Carden, Peter A. Hancock. 2020. “Putting the humanity into inhuman systems: How human factors and ergonomics can be used to manage the risks associated with artificial general intelligence”. https://doi.org/10.1002/hfm.20883.
@@
Department for Science, Innovation & Technology (DSIT). 2025. “Future risks of frontier AI (Annex A)”. https://www.gov.uk/government/publications/frontier-ai-capabilities-and-risks-discussion-paper/future-risks-of-frontier-ai-annex-a.
@@
Bernardi, Jamie, Gabriel Mukobi, Hilary Greaves, Lennart Heim, Markus Anderljung. 2025. “Societal Adaptation to Advanced AI”. https://arxiv.org/pdf/2405.10295.
@@
Drexel, Bill, Caleb Withers. 2024. “Catalyzing Crisis: A Primer on Artificial Intelligence, Catastrophes, and National Security.” https://s3.us-east-1.amazonaws.com/files.cnas.org/documents/Catastrophic-AI˙TECH-2024˙Final.pdf.
@@
Bengio, Yoshua, Sören Mindermann, Daniel Privitera, Tamay Besiroglu, Rishi Bommasani, et al. 2025. “International AI Safety Report 2025.” https://internationalaisafetyreport.org/publication/international-ai-safety-report-2025.
@@
Cass-Beggs, Duncan, Stephen Clare, Dawn Dimowo, Zaheed Kara. 2024. “Framework Convention on Global AI Challenges: Accelerating International Cooperation to Ensure Beneficial, Safe and Inclusive AI.” https://www.cigionline.org/static/documents/AI-challenges˙OW6rTMD.pdf.
@@
Bengio, Yoshua, Geoffrey Hinton, Andrew Yao, Dawn Song, Pieter Abbeel, et al. 2024. “Managing Extreme AI Risks amid Rapid Progress.” https://arxiv.org/abs/2310.17688.
@@
Barnett, Peter, Aaron Scher, David Abecassis. 2025. “Technical Requirements for Halting Dangerous AI Activities.” https://arxiv.org/abs/2507.09801.
@@
Park, Peter S., Simon Goldstein, Aidan O’Gara, Michael Chen, Dan Hendrycks. 2023. “AI Deception: A Survey of Examples, Risks, and Potential Solutions.” https://arxiv.org/abs/2308.14752.
@@
Bales, Adam, William D’Alessandro, Cameron Domenico Kirk-Giannini. 2024. “Artificial Intelligence: Arguments for Catastrophic Risk.” https://arxiv.org/abs/2401.15487.
@@
Motwani, Sumeet Ramesh, Mikhail Baranchuk, Martin Strohmeier, Vijay Bolina, Philip H.S. Torr, et al. 2025. “Secret Collusion among AI Agents: Multi-Agent Deception via Steganography.” https://arxiv.org/abs/2402.07510.
@@
Barnett, Peter, Aaron Scher. 2025. “AI Governance to Avoid Extinction: The Strategic Landscape and Actionable Research Questions.” https://arxiv.org/abs/2505.04592.
@@
Jones, Charles I. 2024. “The AI Dilemma: Growth versus Existential Risk.” https://web.stanford.edu/˜chadj/existentialrisk.pdf.
@@
Swoboda, Torben, Risto Uuk, Lode Lauwaert, Andrew P. Rebera, Ann-Katrien Oimann, et al. 2025. “Examining Popular Arguments Against AI Existential Risk: A Philosophical Analysis.” https://arxiv.org/pdf/2501.04064.
@@
Miotti, Andrea. 2025. “Taking control: Policies to address extinction risks from advanced AI.” https://arxiv.org/pdf/2310.20563.
@@
Phuong, Mary, Roland S. Zimmermann, Ziyue Wang, David Lindner, Victoria Krakovna, et al. 2025. “Evaluating Frontier Models for Stealth and Situational Awareness.” https://arxiv.org/pdf/2505.01420.
@@
Bucknall, Benjamin S., Shiri Dori-Hacohen. 2022. “Current and Near-Term AI as a Potential Existential Risk Factor.” https://arxiv.org/abs/2209.10604.
@@
Russell, Stuart. 2022. “Artificial Intelligence and the Problem of Control.” https://link.springer.com/chapter/10.1007/978-3-030-86144-5˙3.
Li, Xiaojian, Haoyuan Shi, Rongwu Xu, Wei Xu. 2025. “AI Awareness.” https://arxiv.org/pdf/2504.20084.
@@
Dung, Leonard. 2025. “Evaluating approaches for reducing catastrophic risks from AI.” https://link.springer.com/article/10.1007/s43681-024-00475-w.
@@
Kingsley, Jeremy, Syedah Ailia Haider, Shreyansh Jain, Adam Green, Jan Copeman, et al. 2024. “AI landscapes: Exploring future scenarios of AI through to 2030.” https://securesustain.org/wp-content/uploads/2024/05/AI-Landscapes-Exploring-Future-Scenarios-of-AI-Through-to-2030.pdf.
@@
Lynch, Aengus, Benjamin Wright, Caleb Larson, Stuart J. Ritchie, Kevin K. Troy, et al. 2025. “Agentic Misalignment: How LLMs could be insider threats.” https://www.anthropic.com/research/agentic-misalignment.
@@
Benavoli, Alessio, Alessandro Facchini, Marco Zaffalon. 2025. “The AI off-switch problem as a signalling game: bounded rationality and incomparability.” https://arxiv.org/abs/2502.06403.